OpenStudy (anonymous):

If a system does 45 J of work and receives 28 J of heat, what is the Delta E for this change?

OpenStudy (anonymous):

DE = q + w q = 28J (+ indicates received by the system) w = -45J (- indicates word done by the system) DE = (28J) +(-45J) = -17J Peace <3

OpenStudy (anonymous):

thank you! <3